wer [18A, 3C) If f(x) = -3x + 5 and g(x) = 9x – 2, what is the slope of the graph of the function (f-g)(x)?
Answer #1

Given f(x) -3x+5 g(x) = 92-2 Now (f-g)(x) = f(x)-g(x) = €3x+S)-(92-2) =-1200 +7 Slope of the graph = = d 6-9)(x) ef1a2+7) =-